# Proficient Graph Neural Network Design by Accumulating Knowledge on Large Language Models

### v2

@@ Abstract
High-level automation is increasingly critical in AI, driven by rapid advances in large language models (LLMs) and AI agents. However, LLMs, despite their general reasoning power, struggle significantly in specialized, data-sensitive tasks such as designing Graph Neural Networks (GNNs). This difficulty arises from (1) the inherent knowledge gaps in modeling the intricate, varying relationships between graph properties and suitable architectures and (2) the external noise from misleading descriptive inputs, often resulting in generic or even misleading model suggestions. Achieving proficiency in designing data-aware models—defined as the meta-level capability to systematically accumulate, interpret, and apply data-specific design knowledge—remains challenging for existing automated approaches, due to their inefficient construction and application of meta-knowledge. To achieve the meta-level proficiency, we propose DesiGNN, a knowledge-centered framework that systematically converts past model design experiences into structured, fine-grained knowledge priors well fitted to meta-learning with LLMs. To account for the inherent variability and external noise, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs. By constructing a solid meta-knowledge between unseen graph understanding and known effective architecture patterns, DesiGNN can deliver top-5.77% initial model proposals for unseen datasets within seconds, and achieve consistently superior performance with minimal search costs against baselines.
1 Introduction
Automation has become a defining trend in artificial intelligence, reshaping how complex tasks are executed at scale without human intervention or the requirement of deep expertise. From code generation [7, 9, 60] and agentic workflow orchestration [57, 30, 64] to the multi-task generalization of foundation models [29], this paradigm shift is transforming how expertise is embedded and operationalized in AI systems. Among various frontiers, the automatic design of machine learning models stands as one of the most fundamental and impactful challenges in this automation landscape. In the past, automated machine learning (AutoML) [44, 45, 48, 62] has over-emphasized model search specific to independent data and tasks from scratch, which does not conform to the cross-domain and cross-task automation trends. The recent breakthroughs in large language models (LLMs) [11, 4, 41] and AI agents [25, 19, 63] with strong academic backgrounds and contextual reasoning have further fueled expectations—could they autonomously generate tailored models for unseen tasks, thereby democratizing machine learning?
Inspired by LLMs’ general reasoning and generation capability, recent studies [40] have leveraged LLMs to replace human involvement beyond traditional abilities of AutoML, such as task understanding [43, 12] and problem formulation [49]. Despite democratizing complex processes of designing models, realizing highly data-aware automated model design, especially for data-sensitive domains like graph learning, remains elusive for LLMs. Unlike tasks in vision or language domains where LLMs excel in proposing generic solutions that could dominate all tasks and data properties, designing Graph Neural Networks (GNNs) [24, 59, 52, 10, 18, 42] for specific graph properties and graph learning tasks requires a nuanced understanding of how data topology influences model architecture [54, 47]. For example, the homophily nature behind GCN’s aggregation has catastrophic effects when propagating information over heterophily graphs, where two nodes with different labels are more likely to be linked. Unfortunately, current LLM-based methods often fully rely on LLMs’ inherent knowledge and reasoning ability, lacking fine-grained design knowledge about the relationship between graph-specific properties and effective GNN configurations, while also being sensitive to external input noise. As observed in Figure 1(a), the LLM agents adopted in recent methods [43, 12, 49] tend to suggest generally well-performing yet overly generic architectures, inherently falling short in tailoring designs to unseen graph structures. Worse still, as shown in Figure 1(b), they are vulnerable to “artificial hallucinations”—being easily misled by external, irrelevant user descriptions (e.g., being a citation network) or blindly enumerated data statistics that do not accurately reflect the data-aware challenges in model design, resulting in misleading knowledge association and ineffective designs.
@@ 1 Introduction
Figure 1: Case studies illustrating inherent and external limitations of existing LLM-based methods.
@@
Achieving proficiency in designing data-aware models thus demands far more than general LLM-based model design methods, which calls for a meta-level capability to systematically accumulate, interpret, and apply fine-grained, data-specific design knowledge. Humans, akin to training a new Ph.D. student, face substantial hurdles in accumulating this complex knowledge, often undergoing a lengthy learning process to design tailored models proficiently. Meta-based AutoML methods represent initial efforts toward this goal computationally, seeking transferable insights across tasks via learned task embeddings [54, 1, 26] or surrogate predictors [50, 51]. These methods, however, often rely on simple, predefined meta-features that may fail to generalize robustly when facing complex data diversity, thus bending the current automation trend. Extending this concept to graphs, AutoTransfer [5] has made the first promising strides by creating task-model repositories and embedding-based knowledge transfers to improve search efficiency. Nonetheless, it remains limited by predefined task embeddings and the rigid aggregation of prior tasks into random search, insufficiently capturing the fine-grained, graph-aware meta-knowledge documented in the vast graph learning literature [54, 47].
@@
To overcome these limitations, we propose DesiGNN, a knowledge-centered and proficiency-oriented framework for the automated design of data-aware GNNs for unseen tasks. DesiGNN systematically converts past model design experiences into structured, fine-grained knowledge priors that are well fitted to meta-learning with LLMs, novelly bridging the long-standing gap between unseen graph understanding and known effective architecture patterns in graph application research. To address the observed external noise from descriptive inputs and the inherent variability in graph properties, DesiGNN constructs an empirically validated meta-level understanding about graph-GNN-performance from benchmarks and adaptively aligns key properties with graph-specific insights elicited from the graph learning literature, using LLMs as knowledge aligners to enhance the constructed meta-knowledge. This hybridization of empirical knowledge and LLM-guided reasoning empowers DesiGNN to deliver immediate, data-aware model proposals without any training, efficiently refine these architectures based on accumulated insights, and ultimately deliver superior performance with minimal computational overhead. Our contributions are as follows:
@@
  • We empirically identify two key limitations in current LLM-based GNN design approaches: the detrimental external noise from descriptive inputs and the inherent gap in fine-grained design knowledge. Motivated by these insights, we propose a knowledge-centered framework that enhances meta-level proficiency in designing tailored, data-aware GNNs for unseen tasks by systematically accumulating, interpreting, and applying fine-grained meta-knowledge.
@@
  • To account for the inherent variability of graph properties to GNN performance, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs, constructing a solid meta-knowledge between graph topological properties and high-performing GNN architectures. This enables DesiGNN to proficiently deliver immediate, data-aware GNN recommendations without any direct training on the unseen datasets.
@@
  • Extensive experiments across 8 benchmarks and 3 out-of-distribution datasets demonstrate that DesiGNN delivers initial GNN proposals ranking in the top 5.77% of all possible architectures within seconds. Moreover, its knowledge-driven refinement strategy rapidly optimizes these initial suggestions, consistently outperforming automated methods in both performance and efficiency.
2 Related Work
2.1 Automated GNNs
Automated GNNs (AutoGNNs) [44, 15, 62] aim to automatically find an optimal GNN architecture $\theta$ for an unseen graph $G^{u}$ using a controller $\pi$ parameterized by $\alpha$, where the architecture $\theta$ with network weight $\omega$ can achieve best performance $\mathcal{H}(\theta,\omega;G^{u})$ on $G^{u}$. In AutoGNNs, the search spaces are categorized into intra-layer [14] and inter-layer [21] designs, with strategies including reinforcement learning [14], evolutionary algorithms [39], and differentiable search [46]. However, as introduced in Section 1, these approaches lack proficiency and demand substantial computational resources [33], as they optimize configurations $\theta$ without leveraging prior knowledge and require extensive sampling from $\pi_{\alpha}(\theta)$. Recently, meta-learning [13, 6] approaches have sought to overcome these limitations by transferring insights from past tasks through learned task embeddings [54, 5, 1] or surrogate predictors [50, 51, 28]. While these methods mark an initial step toward knowledge-driven AutoGNNs, they rely on coarse, predefined meta-features, often incapable of capturing the generalizable relationships between diverse data properties and effective model configurations.
NAS Benchmark for Designing GNNs. NAS-Bench-Graph [35] offers a standardized benchmark for GNN architecture search, comprising nine datasets and 26,206 unique architectures with associated performance records. Similarly, there are many other benchmarks created by pioneers in different domains [8]. While these benchmarks could provide valuable insights into model-performance relationships, existing methods underutilize their potential for knowledge-driven design due to the absence of a data dimension, which is the foundation of knowledge transfer. Therefore, only a few methods [54, 5] attempted to design feasible models for unseen tasks using such model-performance knowledge. Notably, they rely predominantly on semi-supervised or supervised techniques to fit the missing associations with the data, i.e., require direct training on the unseen data.
2.2 LLMs for Designing GNNs and the Limitations
Inspired by LLMs’ ability to solve optimization tasks [53], recent studies explore LLMs for automating model design [40, 61, 22, 60, 55, 58, 17], focusing on managing task formulation, architecture generation, and optimization. Auto2Graph [49] democratizes the usage of AutoGNNs by translating descriptive user inputs (e.g., Cora [37] is a citation network) into configuration specifications, while GPT4GNAS [43] and GHGNAS [12] iteratively refine designs through crafted prompts, focusing on different graph geneity. Despite democratizing the usage, they suffer from: (1) Over-reliance on external descriptive user inputs that result in poor alignment between graph properties and architecture designs; (2) Raw LLMs inherently lack deep expertise in fine-grained graph-GNN-performance relationships, limiting their ability to suggest data-aware models; (3) Refining model designs solely based on feedback from testing models, lacking actionable, knowledge-driven search strategies.
3 DesiGNN: Proficient Graph Neural Network Design
@@
Figure 2: The illustration of our knowledge-centered DesiGNN pipeline for designing GNNs.
@@
As discussed in Section 1, current LLM-based methods suffer from the lack of proficiency in designing data-aware GNNs, where inherent knowledge gaps and external noise are two major barriers. To formalize these challenges, we adopt a Bayesian perspective to represent the automated GNNs design as solving the posterior belief over a candidate architecture $\theta_{i}$ when facing an unseen graph $G^{u}$:
$\mathbb{P}(\theta_{i}|G^{u})=\frac{\mathbb{P}(G^{u}|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(G^{u})},$ (1)
@@
where $\mathbb{P}(\theta_{i})$ is derived from a knowledge pool $\mathcal{K}$ accumulated from prior tasks $\{G^{i}\}$. However, it remains a non-trivial problem to solve pseudo-likelihood $\mathbb{P}(G^{u}|\theta_{i})$ without formally testing the model on unseen data—a common proficiency bottleneck in existing AutoML that leads to exhaustive search. This proficiency denotes the capability to systematically accumulate, interpret, and apply fine-grained, data-aware meta-knowledge relating graph properties to effective GNN architectures. From a Bayesian perspective, such capability can be conceptually represented as solving the posterior belief $\mathbb{P}(\theta_{i}|G^{u})$ with minimal or even without testing candidate models $\{\theta_{i}\}$ on $G^{u}$.
@@
To enhance such meta-level proficiency without directly testing the unseen data, DesiGNN explicitly integrates empirically validated graph-property knowledge with adaptive LLM alignment to construct a trustworthy meta-level understanding $\mathcal{M}$ between graphs $\mathbf{G}$, GNNs $\mathbf{\Theta}$, and performances $\mathbf{P}$: $\mathcal{M}:(\mathbf{G},\mathbf{\Theta})\rightsquigarrow\mathbf{P}$ based on crucial insights from graph literature [54, 47]. Figure 2 illustrates our Bayesian-inspired pipeline: Graph Understanding, Knowledge Retrieval, and Model Suggestion and Refinement, each component systematically addressing external noise of descriptive data inputs and inherent meta-knowledge gaps, followed by how to strategically apply the constructed meta-knowledge in designing data-aware models. We present our pseudo-code in Algorithm 1 in Appendix A.1.
3.1 Connect Graph Topology to GNN Performance
@@
(b) Descriptive inputs induce invalid task associations that deviate from empirical truth. DesiGNN corr.=0.52.
@@
Figure 3: Empirical studies of two barriers: inherent property variability and external noise.

### v3

@@ Abstract.
High-level automation is increasingly critical in AI, driven by rapid advances in large language models (LLMs) and AI agents. However, LLMs, despite their general reasoning power, struggle significantly in specialized, data-sensitive tasks such as designing Graph Neural Networks (GNNs). This difficulty arises from (1) the inherent knowledge gaps in modeling the intricate, varying relationships between graph properties and suitable architectures and (2) the external noise from misleading descriptive inputs, often resulting in generic or even misleading model suggestions. Achieving proficiency in designing data-aware models—defined as the meta-level capability to systematically accumulate, interpret, and apply data-specific design knowledge—remains challenging for existing automated approaches, due to their inefficient construction and application of meta-knowledge. To achieve meta-level proficiency, we propose DesiGNN, a knowledge-centered framework that systematically converts past model design experience into structured, fine-grained knowledge priors well-suited for meta-learning with LLMs. To account for the inherent variability and external noise, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs. By constructing a solid meta-knowledge between unseen graph understanding and known effective architecture patterns, DesiGNN can deliver top-5.77% initial model proposals for unseen datasets within seconds and achieve consistently superior performance with minimal search cost compared to baselines.
Neural Architecture Search; Graph Neural Networks; Large Language Models
journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USA.booktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAisbn: 979-8-4007-2292-9/2026/02doi: 10.1145/3773966.3777982ccs: Computing methodologies Search methodologiesccs: Computing methodologies Neural networksccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
@@
Figure 1. Case studies illustrating inherent and external limitations of existing LLM-based methods.
Automation has become a defining trend in artificial intelligence, reshaping how complex tasks are executed at scale without human intervention or the requirement of deep expertise. From code generation (Cheng et al., 2023; Choi et al., 2023; Zhang et al., 2023b) and agentic workflow orchestration (Zhang et al., 2024; Liu et al., 2023; Zhuge et al., 2024; Xu et al., 2025; Ouyang et al., 2025) to the multi-task generalization of foundation models (Liu et al., 2025), this paradigm shift is transforming how expertise is embedded and operationalized in AI systems. Among various frontiers, the automated design of machine learning models stands as one of the most fundamental and impactful challenges in this automation landscape. In the past, automated machine learning (AutoML) (Wang et al., 2022a, b; Wei et al., 2021; Zhou et al., 2022) has over-emphasized model search specific to independent data and tasks from scratch, which does not conform to the cross-domain and cross-task automation trends. The recent breakthroughs in large language models (LLMs) (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023; Yue et al., 2025a) and AI agents (Krishnan, 2025; Zhu et al., 2025; Yue et al., 2025b) with strong academic backgrounds and contextual reasoning have further fueled expectations—could they autonomously generate tailored models for unseen tasks, thereby democratizing machine learning?
Inspired by LLMs’ general reasoning and generation capabilities, recent studies (Tornede et al., 2023) have leveraged LLMs to replace human involvement beyond traditional abilities of AutoML, such as task understanding (Wang et al., 2023a; Dong et al., 2023) and problem formulation (Wei et al., 2023). Despite democratizing the complex process of model design, realizing highly data-aware automated model design, especially in data-sensitive domains like graph learning, remains elusive for LLMs. Unlike tasks in vision or language domains, where LLMs excel in proposing generic solutions that could dominate all tasks and data properties, designing Graph Neural Networks (GNNs) (Kipf and Welling, 2016; Zhang and Chen, 2018; Xu et al., 2018; Defferrard et al., 2016; Hamilton et al., 2017; Veličković et al., 2017) for specific graph properties and graph learning tasks requires a nuanced understanding of how data topology influences model architecture (You et al., 2020; Wang et al., 2023b). For example, the homophily nature behind GCN’s aggregation has catastrophic effects when propagating information over heterophily graphs, where two nodes with different labels are more likely to be linked. Unfortunately, current LLM-based methods often rely entirely on LLMs’ inherent knowledge and reasoning, lacking fine-grained design knowledge of the relationships between graph-specific properties and effective GNN configurations, and are also sensitive to external input noise. As observed in Figure 1(a), the LLM agents adopted in recent methods (Wang et al., 2023a; Dong et al., 2023; Wei et al., 2023) tend to suggest architectures that are generally well-performing yet overly generic, inherently falling short in tailoring designs to unseen graph structures. Worse still, as shown in Figure 1(b), they are vulnerable to “artificial hallucinations”—being easily misled by external, irrelevant user descriptions (e.g., being a citation network) or blindly enumerated data statistics that do not accurately reflect the data-aware challenges in model design, resulting in misleading knowledge association and ineffective designs.
@@
Achieving proficiency in designing data-aware models thus demands far more than general LLM-based model design methods, which call for a meta-level capability to systematically accumulate, interpret, and apply fine-grained, data-specific design knowledge. Humans, akin to training a new Ph.D. student, face substantial hurdles in accumulating this complex knowledge, often undergoing a lengthy learning process to design tailored models proficiently. Meta-based AutoML methods represent initial efforts toward this goal computationally, seeking transferable insights across tasks via learned task embeddings (You et al., 2020; Achille et al., 2019; Le et al., 2022) or surrogate predictors (Wen et al., 2020; White et al., 2021). These methods, however, often rely on simple, predefined meta-features that may fail to generalize robustly to complex data diversity, thus bending the current automation trend. Extending this concept to graphs, AutoTransfer (Cao et al., 2023) has made the first promising strides by creating task-model repositories and embedding-based knowledge transfers to improve search efficiency. Nonetheless, it remains limited by predefined task embeddings and the rigid aggregation of prior tasks into random search, insufficiently capturing the fine-grained, graph-aware meta-knowledge documented in the vast graph learning literature (You et al., 2020; Wang et al., 2023b).
@@
To address these limitations, we propose DesiGNN, a knowledge-centered and proficiency-oriented framework for the automated design of data-aware GNNs for unseen tasks. DesiGNN systematically converts past model design experiences into structured, fine-grained knowledge priors well suited to meta-learning with LLMs, thereby novelly bridging the long-standing gap between unseen graph understanding and known effective architecture patterns in graph application research. To address the observed external noise from descriptive inputs and the inherent variability in graph properties, DesiGNN constructs an empirically validated meta-level understanding of graph-GNN-performance from benchmarks and adaptively aligns key properties with graph-specific insights elicited from the graph learning literature, using LLMs as knowledge aligners to enhance the constructed meta-knowledge. This hybridization of empirical knowledge and LLM-guided reasoning empowers DesiGNN to deliver immediate, data-aware model proposals without any training, efficiently refine these architectures based on accumulated insights, and ultimately deliver superior performance with minimal computational overhead. Our contributions are as follows:
@@
  • We empirically identify two key limitations in current LLM-based GNN design approaches: the external noise from descriptive inputs and the inherent gap in fine-grained design knowledge. Motivated by these insights, we propose a knowledge-centered framework that enhances meta-level proficiency in designing tailored, data-aware GNNs for unseen tasks by systematically accumulating, interpreting, and applying fine-grained meta-knowledge.
@@
  • To account for the inherent variability of graph properties on GNN performance, DesiGNN aligns empirical property filtering from extensive benchmarks with adaptive elicitation of literature insights via LLMs, building a solid meta-knowledge between graph topology and high-performing GNN architectures. This enables DesiGNN to proficiently deliver immediate, data-aware GNN recommendations without any training on the unseen data.
@@
  • Extensive experiments across 3 out-of-distribution datasets and 8 benchmarks demonstrate that DesiGNN delivers initial GNN proposals ranking in the top 5.77% of all possible architectures within seconds. Moreover, its knowledge-driven refinement strategy rapidly optimizes these initial suggestions, consistently outperforming automated methods in both performance and efficiency.
2. Related Work
@@
Figure 2. The illustration of our knowledge-centered DesiGNN pipeline for designing GNNs. The orange module is Graph understanding, the green module is Knowledge Retrieval, and the purple module is Model Suggestion and Refinement.
2.1. Automated GNNs
Automated GNNs (AutoGNNs) (Wang et al., 2022a; Gao et al., 2021; Zhou et al., 2022) aim to automatically find an optimal GNN architecture $\theta$ for an unseen graph $G^{u}$ using a controller $\pi$ parameterized by $\alpha$, where the architecture $\theta$ with network weight $\omega$ can achieve best performance $\mathcal{H}(\theta,\omega;G^{u})$ on $G^{u}$. In AutoGNNs, the search spaces are categorized into intra-layer (Gao et al., 2019) and inter-layer (Huan et al., 2021) designs, with strategies including reinforcement learning (Gao et al., 2019), evolutionary algorithms (Shi et al., 2022), differentiable search (Wang et al., 2021), and search-based model fine-tuning (Zhili et al., 2024). Similar strategies exist for the knowledge graph (Di and Chen, 2023; Shimin et al., 2021; Di et al., 2021; Iyer et al., 2022). However, these approaches lack proficiency and demand substantial computational resources (Oloulade et al., 2021), as they optimize configurations $\theta$ without leveraging prior knowledge and require extensive sampling from $\pi_{\alpha}(\theta)$. Recently, meta-learning (Finn et al., 2017; Chen et al., 2021) methods have sought to overcome these limitations by transferring insights from past tasks through learned task embeddings (Cao et al., 2023; Achille et al., 2019) or surrogate predictors (Wen et al., 2020; White et al., 2021; Li et al., 2023). While these methods mark an initial step toward knowledge-driven AutoGNNs, they rely on coarse, predefined meta-features that are often incapable of capturing generalizable relationships between diverse data properties and effective model configurations.
NAS Benchmark for Designing GNNs. NAS-Bench-Graph (Qin et al., 2022) offers a standard benchmark for GNN architecture search, comprising nine datasets and 26,206 unique architectures with associated performance records. Similarly, many other benchmarks have been created by pioneers across different domains (Chitty-Venkata et al., 2023). While these benchmarks could provide valuable insights into model-performance relationships, existing works underutilize their potential for knowledge-driven design because they lack a data dimension, which is the foundation of knowledge transfer. Therefore, only a few methods (Liu et al., ; You et al., 2020; Cao et al., 2023) attempted to design feasible models for unseen tasks using such model-performance knowledge. Notably, they rely primarily on semi-supervised or supervised techniques to fill in missing data associations, i.e., require direct training on unseen data.
2.2. LLMs for Designing GNNs and Limitations
Inspired by LLMs’ ability to solve optimization tasks (Yang et al., 2023), recent studies explore their use to automate model design (Zheng et al., 2023; Jawahar et al., 2023; Zhang et al., 2023b; Yu et al., 2023; Zhang et al., 2023a; Guo et al., 2024), focusing on managing task formulation, architecture generation, and optimization. Auto2Graph (Wei et al., 2023) democratizes the use of AutoGNNs by translating descriptive user inputs (e.g., Cora (Sen et al., 2008) is a citation network) into configuration specifications. GPT4GNAS (Wang et al., 2023a) and GHGNAS (Dong et al., 2023) iteratively refine model designs through crafted prompts, focusing on different graph geneity. Despite democratizing the usage, they suffer from: (1) Over-reliance on external descriptive user inputs that result in poor alignment between graph properties and architecture designs; (2) Raw LLMs inherently lack deep expertise in fine-grained graph-GNN-performance relationships, limiting their ability to suggest data-aware models; (3) Refining model designs solely based on feedback from testing models, lacking actionable, knowledge-driven search strategies.
3. Proficient Graph Neural Network Design
@@
As discussed in Section 1, current LLM-based methods lack proficiency in designing data-aware GNNs, with inherent knowledge gaps and external noise as major barriers. To formalize these challenges, we adopt a Bayesian perspective to represent the automated GNNs design as solving the posterior belief over candidate architecture $\theta_{i}$ when facing an unseen graph $G^{u}$:
(1) $\mathbb{P}(\theta_{i}|G^{u})=\frac{\mathbb{P}(G^{u}|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(G^{u})},$
@@
where $\mathbb{P}(\theta_{i})$ is derived from a knowledge pool $\mathcal{K}$ accumulated from prior tasks $\{G^{i}\}$. However, it remains a non-trivial problem to solve pseudo-likelihood $\mathbb{P}(G^{u}|\theta_{i})$ without formally testing the model on unseen data—a common proficiency bottleneck in existing AutoML that leads to exhaustive search. This proficiency denotes the capability to systematically accumulate, interpret, and apply fine-grained, data-aware meta-knowledge relating graph properties to effective GNN architectures. In our formulation, such capability can be conceptually represented as solving for the posterior belief $\mathbb{P}(\theta_{i}|G^{u})$, with minimal or even no testing of candidate models $\{\theta_{i}\}$ on $G^{u}$.
@@
To enhance such meta-level proficiency without directly testing the unseen data, DesiGNN explicitly integrates empirically validated graph-property knowledge with adaptive LLM alignment to construct a trustworthy meta-level understanding $\mathcal{M}$ between graphs $\mathbf{G}$, GNNs $\mathbf{\Theta}$, and performances $\mathbf{P}$: $\mathcal{M}:(\mathbf{G},\mathbf{\Theta})\rightsquigarrow\mathbf{P}$ based on crucial insights from graph literature (You et al., 2020; Wang et al., 2023b). Figure 2 illustrates our Bayesian-inspired pipeline: Graph Understanding, Knowledge Retrieval, and Model Suggestion and Refinement, each component systematically addressing external noise in descriptive data inputs and inherent meta-knowledge gaps, and how to strategically apply the constructed meta-knowledge in designing data-aware models.
@@
(b) Descriptive inputs induce invalid task associations that deviate from empirical truth. DesiGNN exhibits a better correlation of 0.52.
@@
Figure 3. Empirical studies of two barriers: inherent property variability and external noise.
3.1. Connect Graph Topology to GNN Design